### v2

# Oriented matroids and type $\mathbb{A}$ cluster categories

Abstract.
For any cluster-tilting object $\mathsf{T}$ in the cluster category $\mathscr{C}_{n}$ of type $\mathbb{A}_{n}$, we construct a rank-four oriented matroid $\mathcal{M}_{\mathsf{T}}$ such that stackable triangulations of $\mathcal{M}_{\mathsf{T}}$ are in bijection with equivalence classes of maximal green sequences with initial cluster $\mathsf{T}$. This generalises the result that equivalence classes of maximal green sequences of linearly oriented $\mathbb{A}_{n}$ are in bijection with triangulations of a three-dimensional cyclic polytope. The definition of the oriented matroid $\mathcal{M}_{\mathsf{T}}$ arises from the extriangulated structure on $\mathscr{C}_{n}$ which makes $\mathsf{T}$ projective.
Key words and phrases:
Oriented matroids, triangulations, maximal green sequences, cluster categories
2020 Mathematics Subject Classification:
Primary: 16G20; Secondary: 52C40, 13F60
I would like to thank Rudradip Biswas for the opportunity to speak about a preliminary version of this work at the “New Directions in Group Theory and Triangulated Categories” seminar, and Mikhail Gorsky for the reference [8]. I am currently supported by EPSRC grant EP/W001780/1.
1. Introduction
From the very beginning, there has been an intimate relationship between cluster algebras and the combinatorics of triangulations. The most fundamental example of this is the $\mathbb{A}_{n}$ cluster algebra [10], whose cluster variables are in bijection with arcs in a convex $(n+3)$-gon, with clusters in bijection with triangulations, and mutation corresponding to flipping an arc in a quadrilateral. This was extended to cluster algebras from triangulations of surfaces in [9].
Higher-even-dimensional triangulations were connected with cluster algebras in [16], and higher odd dimensions in [22]. In particular, in [22] it was shown that equivalence classes of maximal green sequences of a linearly oriented type $\mathbb{A}$ quiver were in bijection with triangulations of a three-dimensional cyclic polytope. Maximal green sequences are finite sequences of mutations in a cluster algebra, first introduced into the mathematics literature by Keller [13], and into the physics literature by Cecotti, Cordova, and Vafa [6, Section 4.3]. In fact, these physicists already made a connection between maximal green sequences and three-dimensional triangulations [6, Section 5]. This appearance of maximal green sequences in physics is related to their role in quantum dilogarithm identities [20] and the Kontsevich–Soibelman wall-crossing formula from Donaldson–Thomas theory [14].
Let us explain maximal green sequences in more detail. Choosing an initial cluster in a cluster algebra orients mutations so that some are forwards and some are backwards. There are combinatorial rules for colouring the vertices of a quiver green or red such that forwards mutation corresponds to mutating at a green vertex, which then turns it red; backwards mutation is then the reverse of this. In this picture, a maximal green sequence is a finite sequence of mutations at green vertices such that the quiver turns from all green to all red. Forwards mutation may thus be called “green mutation”, with backwards mutation called “red”.
Given the bijection from [22] between triangulations of three-dimensional cyclic polytopes and equivalence classes of maximal green sequences of linearly-oriented $\mathbb{A}_{n}$, the following natural question then arises. Given an arbitrary orientation of the type $\mathbb{A}$ Dynkin diagram, is there a three-dimensional polytope whose triangulations are in bijection with its maximal green sequences, up to equivalence? Following [22], this bijection should be such that the simplices in the polytope are in bijection with the mutations in the maximal green sequence, so that the number of simplices of the triangulation equals the length of the maximal green sequence.
One can deduce what properties such a three-dimensional polytope should have. An exchange pair in a maximal green sequence corresponds to flipping an arc in a quadrilateral. As in [22], one should lift this quadrilateral to a $3$-simplex such that the green mutation goes from the lower side of the simplex to the upper side. This amounts to choosing an orientation for each $3$-simplex in our three-dimensional polytope, and this determines the corresponding oriented matroid.
In order to determine the orientation of each $3$-simplex, we need a criterion which tells us, given an initial cluster, which mutations are green. Such a criterion is given by [8] in terms of the cluster category of [4], where a choice of initial cluster is given by a choice of cluster-tilting object $\mathsf{T}$. In particular, different orientations of $\mathbb{A}_{n}$ can be achieved by different choices of $\mathsf{T}$. Of course, it is not obvious that orienting the $3$-simplices according to which exchange pairs are made green by the cluster-tilting object $\mathsf{T}$ gives a well-defined oriented matroid $\mathcal{M}_{\mathsf{T}}$: certain axioms need to be satisfied. The first main theorem of this paper is that these axioms do in fact hold. We moreover conjecture that $\mathcal{M}_{\mathsf{T}}$ is indeed the oriented matroid of some polytope: that $\mathcal{M}_{\mathsf{T}}$ is “realisable”.
Theorem A (Theorem 3.8).
Given a cluster-tilting object $\mathsf{T}$ in the cluster category $\mathscr{C}_{n}$ of type $\mathbb{A}_{n}$, we have that $\mathcal{M}_{\mathsf{T}}$ is a well-defined oriented matroid.
The oriented matroids $\mathcal{M}_{\mathsf{T}}$ are orientations of the uniform matroid: all possible $3$-simplices are bases of the matroid. Orientations of the uniform matroid have been studied recently in the context of their associated “chirotropical Grassmannians”, which were introduced in [5] in the context of scattering amplitudes, and subsequently studied in [1]. While uniform matroids are always realisable [2, Section 8.3], uniform oriented matroids are not always realisable [2, Proposition 8.3.2].
Having proved that the oriented matroids $\mathcal{M}_{\mathsf{T}}$ are well-defined, we construct the desired bijection between equivalence classes of maximal green sequences with initial cluster $\mathsf{T}$ and the subset of triangulations of $\mathcal{M}_{\mathsf{T}}$ which are stackable.
Theorem B (Theorem 4.11).
Let $\mathsf{T}$ be a basic cluster-tilting object in $\mathscr{C}_{n}$. There is then a bijection between stackable triangulations of $\mathcal{M}_{\mathsf{T}}$ and equivalence classes of maximal green sequences with initial cluster $\mathsf{T}$.
We conjecture that in fact all triangulations of $\mathcal{M}_{\mathsf{T}}$ are stackable. If this is so, and $\mathcal{M}_{\mathsf{T}}$ is realisable, then for every basic cluster-tilting object $\mathsf{T}\in\mathscr{C}_{n}$, there is a polytope whose triangulations are in bijection with equivalence classes of maximal green sequences with initial cluster $\mathsf{T}$.
This paper is structured as follows. In Section 2 we give requisite background on oriented matroids, their triangulations, cluster categories, and maximal green sequences. In Section 3, we construct the oriented matroids and prove Theorem A. In Section 4, we then prove the relation between their triangulations and maximal green sequences, and prove Theorem B.
2. Background
2.1. Oriented matroids
A matroid is a structure abstracting the linear dependence relations that hold among a finite set of vectors. An oriented matroid is a matroid with some extra structure — an “orientation” — which also keeps track of the signs of the coefficients in these linear relations. We will follow the standard reference for oriented matroids [2].
Matroids are famous for having many different “cryptomorphic” definitions. One such definition axiomatises the notion of the “bases” of a matroid, which correspond to bases in the standard sense of linear algebra. Using the bases definition of a matroid, an orientation of a matroid is given by assigning a sign to every ordered basis, corresponding to the sign of the determinant of the ordered basis. This assignment of signs gives an object called a chirotope, which we now describe.
Definition 2.1 ([2, Definition 3.5.3]).
Let $r\geqslant 1$ be an integer, and let $E$ be a finite set, known as the ground set. A chirotope of rank $r$ on $E$ is a map $\chi\colon E^{r}\to\{-1,0,+1\}$ which satisfies the following three properties.
  1. (1)
    $\chi$ is not identically zero.
  2. (2)
    $\chi$ is alternating, that is,
    $\chi(x_{\pi(1)},x_{\pi(2)},\dots,x_{\pi(r)})=\mathsf{sign}(\pi)(x_{1},x_{2},\dots,x_{r})$
    for all $x_{1},x_{2},\dots,x_{r}\in E$ and every permutation $\pi$ of $r]:=\{1,2,\dots,r\}$.
  3. (3)
    For all $x_{1},x_{2},\dots,x_{r},y_{1},y_{2},\dots,y_{r}\in E$ such that
    $\chi(y_{i},x_{2},x_{3},\dots,x_{r})\chi(y_{1},y_{2},\dots,y_{i-1},x_{1},y_{i+1},y_{i+2},\dots,y_{r})\geqslant 0$
    for all $i\in[r]$, we have
    $\chi(x_{1},x_{2},\dots,x_{r})\chi(y_{1},y_{2},\dots,y_{r})\geqslant 0.$
The chirotope $\chi$ is realisable if there exists a map $\rho\colon E\to\mathbb{R}^{r}$ such that
$\chi(x_{1},x_{2},\dots,x_{r})=\mathsf{sign}\,\mathsf{Det}\,(\rho(x_{1})\rho(x_{2})\dots\rho(x_{r})),$
where here we have specified a matrix by its column vectors.
Aside from the definition of matroids using bases, one can also define them using “circuits”. In terms of linear algebra, circuits are minimal linearly dependent sets. That is, a circuit is a linearly dependent set whose subsets are all linearly independent. Oriented matroids have signed circuits, which are signed sets.
Definition 2.2 ([2, p.102]).
A signed set $X$ is a set $\underline{X}$ together with a partition $(X^{+},X^{-})$ of $\underline{X}$ into two distinguished subsets. Here $\underline{X}$ is called the support of $X$. Given a set $E$, a signed subset of $E$ is a signed set $X$ with $\underline{X}\subseteq E$.
One can specify axioms for a set of signed subsets of the ground set $E$ to be the set of signed circuits of an oriented matroid [2, Definition 3.2.1]. Note that we do not identify an oriented matroid with its chirotope or with its set of signed circuits. We think of oriented matroids as objects in the style of object-oriented programming: an oriented matroid has a chirotope and a set of signed circuits as attributes; it is determined by specifying either of these pieces of data, but it is not identified with either of them. We say that an oriented matroid is realisable if its chirotope is realisable.
In this paper, it will be useful to restrict an oriented matroid to a subset.
Proposition 2.3 ([2, Proposition 3.3.1, pp.133–4]).
Let $\mathcal{M}$ be an oriented matroid on a ground set $E$ with set of signed circuits $\mathcal{C}$ and let $V$ be a subset of $E$. Then $\mathcal{C}(V):=\{C\in\mathcal{C}:\underline{C}\subseteq V\}$, the set of circuits of $\mathcal{M}$ contained in $V$, is the set of signed circuits of an oriented matroid on $V$. We call this oriented matroid the restriction of $\mathcal{M}$ to $V$, and denote it by $\mathcal{M}(V)$.
Let $\chi\colon E^{r}\to\{-1,0,+1\}$ be the chirotope of $\mathcal{M}$. If $\mathcal{M}(V)$ also has rank $r$, then its chirotope is $\chi|_{V}$.
For every oriented matroid $\mathcal{M}$ on a set $E$, there is a dual oriented matroid $\mathcal{M}^{\ast}$ on $E$ [2, Proposition 3.4.1]. The circuits of $\mathcal{M}^{\ast}$ are known as the cocircuits of $\mathcal{M}$. The set of cocircuits of $\mathcal{M}$ is denoted $\mathcal{C}^{\ast}$. If $\mathcal{M}$ is realisable, then cocircuits of $\mathcal{M}$ correspond to hyperplanes, with one half of the cocircuit lying on one side of the hyperplane, the other half lying on the other [2, p.116], and the complement of the cocircuit lying in the hyperplane.
Definition 2.4 ([2, p.377]).
A signed circuit $X$ is positive if $\underline{X}=X^{+}$. An oriented matroid $\mathcal{M}$ is acyclic if it does not have a positive circuit.
Let $\mathcal{M}$ be an acyclic oriented matroid on a set $E$. A hyperplane of $\mathcal{M}$ is a subset $H\subseteq E$ such that $E\setminus H$ is a cocircuit of $\mathcal{M}$. An open halfspace of $\mathcal{M}$ is $Y^{+}$ for $Y$ a cocircuit of $\mathcal{M}$. A facet of $\mathcal{M}$ is a hyperplane $H$ such that $E\setminus H$ is an open halfspace. A face of $\mathcal{M}$ is an intersection of facets of $\mathcal{M}$.
Finally, we define triangulations of oriented matroids.
Definition 2.5 ([21, Theorem 2.4(f)]).
Let $\mathcal{M}$ be an acyclic oriented matroid of rank $r$ on a set $E$. A non-empty collection $\Delta$ of bases of $\mathcal{M}$ is called a triangulation of $\mathcal{M}$ if the following conditions hold.
  1. (1)
    If $\sigma,\tau\in\Delta$, then $\sigma\cap\tau$ is a common face of the restrictions $\mathcal{M}(\sigma)$ and $\mathcal{M}(\tau)$.
  2. (2)
    If $\sigma\in\Delta$, then each facet of $\mathcal{M}(\sigma)$ is either a facet of $\mathcal{M}$ or is contained in precisely two cells of $\Delta$.
  3. (3)
    No two bases $\sigma$ and $\tau$ overlap on a circuit $C=(C^{+},C^{-})$, meaning that $C^{+}\subseteq\sigma$ and there is an element $a\in C^{+}$ such that $\underline{C}\setminus\{a\}\subseteq\tau$.
We also refer to the bases in $\Delta$ as (r1)(r-1)-simplices.
2.2. Cluster categories of type $\mathbb{A}$
Cluster categories were introduced in the paper [4], which we refer to for full background. Given an acyclic quiver $Q$, its path algebra is denoted $KQ$, where $K$ is a field. We write $\mathscr{D}^{b}(KQ)$ for the bounded derived category of right $KQ$-modules which are finite-dimensional over $K$. The cluster category $\mathscr{C}_{Q}$ of $Q$ is the orbit category $\mathscr{D}^{b}(KQ)/\nu\Sigma^{-2}$, where $\nu:=D(KQ)\otimes^{\mathbf{L}}_{KQ}-\colon\mathscr{D}^{b}(KQ)\to\mathscr{D}^{b}(KQ)$ is the Nakayama functor and $\Sigma$ is the suspension of $\mathscr{D}^{b}(KQ)$. Here $D(-)=\operatorname{Hom}_{K}(-,K)$ is the standard duality.
The linearly oriented $\mathbb{A}_{n}$ quiver is $1\to 2\to\dots\to n$ and we denote its cluster category by $\mathscr{C}_{n}$. In fact, $\mathscr{C}_{n}$ is independent of the choice of orientation of the quiver. The cluster category $\mathscr{C}_{n}$ is a $K$-linear triangulated category with finitely many indecomposable objects, and one can describe its morphisms and suspension functor combinatorially. Indeed, up to isomorphism, the indecomposable objects of $\mathscr{C}_{n}$ may be labelled $\mathsf{O}_{B}$ for $B=\{b_{1},b_{2}\}\in\binom{[n+3]}{2}$ with $b_{1}\neq b_{2}\pm 1$. Here, when we write $b_{2}\pm 1$, we are using modular arithmetic with representatives in $[n+3]$, rather than $\{0,1,\dots,n+2\}$, as is more usual; we will continue do this in this paper. Hence, indecomposable objects in $\mathscr{C}_{n}$ are in bijection with arcs in a convex $(n+3)$-gon.
We call a permutation in the subgroup of the symmetric group $\mathfrak{S}_{m}$ generated by the permutation $x\mapsto x+1$ a cyclic permutation. By rotation, we mean applying a cyclic permutation. Given $\{a_{1},a_{2},\dots,a_{l}\}\subseteq[m]$ with $l\geqslant 3$, we write $a_{1}\prec a_{2}\prec\dots\prec a_{l}$ if there is a cyclic permutation $\pi\in\mathfrak{S}_{m}$ such that $\pi(a_{1})<\pi(a_{2})<\dots<\pi(a_{l})$ in the usual ordering on $[m]$. We call the ordering $a_{1}\prec a_{2}\prec\dots\prec a_{l}$ a cyclic ordering. We then have that $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{B},\mathsf{O}_{C})\neq 0$ if and only if $b_{1}-1\prec c_{1}\prec b_{2}-1\prec c_{2}$ with $B=\{b_{1},b_{2}\}$ and $C=\{c_{1},c_{2}\}$. Moreover, we have that if $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{B},\mathsf{O}_{C})\neq 0$, then $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{B},\mathsf{O}_{C})\cong K$. Finally, we have $\Sigma\mathsf{O}_{B}=\mathsf{O}_{B-1}$, where $B\pm 1:=\{b\pm 1:b\in B\}$. These descriptions follow, for instance, from specialising the results of [16].
An object $\mathsf{T}$ of $\mathscr{C}_{n}$ is cluster-tilting if $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{T},\Sigma\mathsf{T})=0$ and it has $n$ isomorphism classes of indecomposable summands. There is a bijection between basic cluster-tilting objects in $\mathscr{C}_{n}$ and triangulations of the $(n+3)$-gon via the map sending a basic cluster-tilting object $\mathsf{T}=\bigoplus_{i=1}^{n}\mathsf{O}_{B_{i}}$ to the triangulation with set of arcs $\{B_{1},B_{2},\dots,B_{n}\}$. This is because the condition of $\mathsf{T}$ being cluster-tilting is equivalent to the condition that none of the arcs $B_{i}$ cross. Recall that basic means that distinct summands are non-isomorphic.
2.3. Maximal green sequences
To suit our purposes, we use a different presentation of maximal green sequences to the usual one. Our presentation uses the concept of an extriangulated category, for background on which, see [15]. Given a cluster-tilting object $\mathsf{T}$ in $\mathscr{C}_{n}$, we consider the extriangulated structure on $\mathscr{C}_{n}$ given by those triangles $\mathsf{X}\to\mathsf{Y}\to\mathsf{Z}\to\Sigma\mathsf{X}$ for which the morphism $\mathsf{Z}\to\Sigma\mathsf{X}$ factors through $\Sigma\mathsf{T}$. We write $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{Z},\mathsf{X})$ for $\operatorname{Ext}$ in this extriangulated structure. This extriangulated structure was first introduced in [17, Section 4.7.1] and is an example of a relative extriangulated structure as studied in [12, Section 3.2].
We say that two basic cluster tilting objects $\mathsf{T}_{1}$ and $\mathsf{T}_{2}$ are related by a mutation if $\mathsf{T}_{1}=\mathsf{E}\oplus\mathsf{X}$ and $\mathsf{T}_{2}=\mathsf{E}\oplus\mathsf{Y}$, where $\mathsf{X}$ and $\mathsf{Y}$ are indecomposable and not isomorphic. We say that the mutation from $\mathsf{T}_{1}$ to $\mathsf{T}_{2}$ is $\mathsf{T}$-green if $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{Y},\mathsf{X})\neq 0$. We then say that $(\mathsf{X},\mathsf{Y})$ is the exchange pair of the mutation.
A maximal green sequence of $\mathsf{T}$ is a sequence $\mathcal{G}=(\mathsf{T}_{0},\mathsf{T}_{1},\dots,\mathsf{T}_{l})$ of basic cluster-tilting objects $\mathsf{T}_{i}\in\mathscr{C}_{n}$ such that $\mathsf{T}_{0}=\mathsf{T}$, $\mathsf{T}_{l}=\Sigma\mathsf{T}$, and for all $i\in[l]$, $\mathsf{T}_{i}$ is a $\mathsf{T}$-green mutation of $\mathsf{T}_{i-1}$. Denote by $(\mathsf{X}_{i},\mathsf{Y}_{i})$ the exchange pair of the $\mathsf{T}$-green mutation from $\mathsf{T}_{i-1}$ to $\mathsf{T}_{i}$. We write $\mathrm{Exch}(\mathcal{G}):=\{(\mathsf{X}_{1},\mathsf{Y}_{1}),(\mathsf{X}_{1},\mathsf{Y}_{1}),\dots,(\mathsf{X}_{l},\mathsf{Y}_{l})\}$ for the set of exchange pairs of $\mathcal{G}$. We say that two maximal green sequences $\mathcal{G}$ and $\mathcal{G}^{\prime}$ of $\mathsf{T}$ are equivalent if $\mathrm{Exch}(\mathcal{G})=\mathrm{Exch}(\mathcal{G}^{\prime})$. See [11, Theorem 3.19] for other characterisations of this equivalence relation.
Remark 2.6.
What we call a maximal green sequence of $\mathsf{T}$ corresponds to a maximal green sequence of $\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T}$ in the literature. We explain how our definition is equivalent to the definition of maximal green sequence using two-term silting complexes (see [11, Section 2.3.5]).
It follows from [8, Theorem 6.15] that the category $K^{[-1,0]}(\mathrm{proj}\,\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T})$ of two-term complexes of projectives over $\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T}$ is equivalent, as an extriangulated category, to the extriangulated ideal quotient of $\mathscr{C}_{n}$ (with the above extriangulated structure) by the morphisms from the shifted projectives to the projectives. It follows from [8, Proposition 5.1] that the relative extriangulated structure used in [8, Theorem 6.15] is the one described above.
The definition of a maximal green sequence of $\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T}$ in terms of two-term silting complexes is a sequence of green mutations of basic two-term silting complexes from the projectives to the shifted projectives. By the equivalence of extriangulated categories from the previous paragraph, this corresponds to our definition of a maximal green sequence of $\mathsf{T}$ as a sequence of $\mathsf{T}$-green mutations from $\mathsf{T}$ to $\Sigma\mathsf{T}$.
3. Constructing the oriented matroids
As in Section 2.2, a cluster-tilting object in $\mathscr{C}_{n}$ corresponds to a triangulation of a convex $(n+3)$-gon. In this section, we will construct an oriented matroid $\mathcal{M}_{T}$ from a triangulation $T$ of a convex $(n+3)$-gon. In Section 4, we will show that if $\Sigma\mathsf{T}$ is the cluster-tilting object in $\mathscr{C}_{n}$ corresponding to $T$, then $\mathcal{M}_{\mathsf{T}}:=\mathcal{M}_{T}$ is the oriented matroid with the desired properties. That is, we will show that stackable triangulations of $\mathcal{M}_{\mathsf{T}}$ are in bijection with equivalence classes of maximal green sequences of $\mathsf{T}$.
For now, let $T$ be a triangulation of a convex $m$-gon, where $m\geqslant 4$. We construct an oriented matroid $\mathcal{M}_{T}$ from $T$ by specifying its chirotope. Given $x,y\in[m]$ with $x\neq y$, we denote the arc with vertices $x$ and $y$ by $xy$. We refer to a subset of $[m]$ of size $k$ as a $k$-subset.
Definition 3.1.
The ground set of the oriented matroid $\mathcal{M}_{T}$ is $[m]$. Let $\{a,b,c,d\}$ be a 4-subset of $[m]$ such that $a\prec b\prec c\prec d$. We then define
[…]
We extend this to tuples which are not cyclically ordered by requiring that $\chi_{T}$ be alternating, in the sense of Definition 2.1(2).
We first need to show that $\chi_{T}$ is well-defined as a map.
Lemma 3.2.
Given $\{a\prec b\prec c\prec d\}\subseteq[m]$, the triangulation $T$ cannot both contain an arc $xy$ with $a\prec b\preccurlyeq x\prec c\prec d\preccurlyeq y$ and an arc $x^{\prime}y^{\prime}$ with $a\preccurlyeq x^{\prime}\prec b\prec c\preccurlyeq y^{\prime}\prec d$.
Proof.
If such arcs $xy$ and $x^{\prime}y^{\prime}$ exist, then we have $a\preccurlyeq x^{\prime}\prec b\preccurlyeq x\prec c\preccurlyeq y^{\prime}\prec d\preccurlyeq y$, and so we have $x^{\prime}\prec x\prec y^{\prime}\prec y$. This means that the arcs $xy$ and $x^{\prime}y^{\prime}$ cross, which contradicts the fact that $T$ is a triangulation. ∎
Lemma 3.3.
Given $\{a\prec b\prec c\prec d\}\subseteq[m]$, the triangulation $T$ either contains an arc $xy$ with $a\prec b\preccurlyeq x\prec c\prec d\preccurlyeq y$ or an arc $xy$ with $a\preccurlyeq x\prec b\prec c\preccurlyeq y\prec d$.
Proof.
If the triangulation $T$ contains arcs $xy$ with $a\preccurlyeq x\prec y\prec b$, then $T$ restricts to a triangulation of the polygon with vertices $V=[m]\setminus\{v\in[m]:x\prec v\prec y\}$, where we also have $V\supseteq\{a,b,c,d\}$. Hence, we can assume that there are no such arcs $xy$ of $T$, and likewise assume that $T$ contains no arcs $xy$ with $b\preccurlyeq x\prec y\prec c$, $c\preccurlyeq x\prec y\prec d$, or $d\preccurlyeq x\prec y\prec a$. Thus, if $x$ is a vertex with $a\preccurlyeq x\prec b$, then there must be an arc (or boundary edge) $xy$ of $T$ with either $b\preccurlyeq y\prec c$, $c\preccurlyeq y\prec d$, or $d\preccurlyeq y\prec a$. If the middle case holds, then we are done, so we can assume ($\ast$) that the first or the third case holds for all $x$ with $a\preccurlyeq x\prec b$. If we have $a\preccurlyeq x\prec x^{\prime}\prec b$, with $x^{\prime}y^{\prime}$ an arc of $T$, then we cannot have $b\preccurlyeq y\prec c$ and $d\preccurlyeq y^{\prime}\prec a$, since then $xy$ and $x^{\prime}y^{\prime}$ cross. Thus, let $\overrightarrow{x}$ be the rightmost vertex with $a\prec\overrightarrow{x}\prec b$ such that there is an arc $\overrightarrow{x}\overrightarrow{y}$ of $T$ with $d\preccurlyeq\overrightarrow{y}\prec a$. Similarly, let $\overleftarrow{x}$ be the leftmost vertex with $a\prec\overleftarrow{x}\prec b$ such that there is an arc $\overleftarrow{x}\overleftarrow{y}$ of $T$ with $b\preccurlyeq\overleftarrow{y}\prec c$. By assumption ($\ast$), there are no vertices $x$ with $\overrightarrow{x}\prec x\prec\overleftarrow{x}$. In fact, one can see that $\overrightarrow{x}=\overleftarrow{x}$, since one can otherwise consider the triangle with the side $\overrightarrow{x}\overleftarrow{x}$, which must either contain an arc $\overrightarrow{x}y$ with $b\preccurlyeq y\prec c$ or an arc $\overleftarrow{x}y$ with $d\preccurlyeq y\prec a$.
We thus let $x_{ab}=\overrightarrow{x}=\overleftarrow{x}$. One can deduce the existence of analogous vertices $x_{bc}$, $x_{cd}$, and $x_{da}$ with $b\preccurlyeq x_{bc}\prec c$, $c\preccurlyeq x_{cd}\prec d$, and $d\preccurlyeq x_{da}\prec a$. It can be seen that $T$ therefore contains a quadrilateral with vertices $x_{ab}$, $x_{bc}$, $x_{cd}$, and $x_{da}$. The diagonal $xy$ of this quadrilateral which is an arc of $T$ then gives us the arc in the statement of the lemma. ∎
We now need to show that $\mathcal{M}_{T}$ is a well-defined oriented matroid. We will use [2, Corollary 3.6.3], which says that $\mathcal{M}_{T}$ is well-defined if every six-element restriction of $\mathcal{M}_{T}$ is realisable. Hence, we want to understand restrictions of $\mathcal{M}_{T}$.
Definition 3.4.
Let $T$ be a triangulation of $[m]$ and let $V\subseteq[m]$ be a subset with $|V|\geqslant 4$. We say that an arc $xy$ of $T$ separates $V$ if $V=V_{x}\sqcup V_{y}$ with $V_{x}$ and $V_{y}$ non-empty, with $V_{x}\preccurlyeq x\prec V_{y}\preccurlyeq y$. By this, we of course mean that for every $a\in V_{x}$ and $c\in V_{y}$, we have $a\preccurlyeq x\prec c\preccurlyeq y$.
Given an arc $xy$ which separates $V$, set $x^{V}:=\max_{\circlearrowleft}V_{x}$ and $y^{V}:=\max_{\circlearrowleft}V_{y}$, where $\max_{\circlearrowleft}V_{x}$ is the unique element of $V_{x}$ such that $v\preccurlyeq\max_{\circlearrowleft}V_{x}\preccurlyeq x$ for all elements $v\in V_{x}$, and $\max_{\circlearrowleft}V_{y}$ is defined similarly. Define $T_{V}:=\{x^{V}y^{V}:xy\text{ an arc of $T$ which separates }V\}$.
Lemma 3.5.
Given a triangulation $T$ of the $m$-gon and $V\subseteq[m]$ with $|V|\geqslant 4$, we have that $T_{V}$ is a triangulation of the convex polygon with vertices $V$.
Proof.
[…]
Proposition 3.6.
Let $T$ be a triangulation of the $m$-gon and $V\subseteq[m]$ with $|V|\geqslant 4$. Then $\chi_{T}|_{V^{4}}=\chi_{T_{V}}$, and so $\mathcal{M}_{T}(V)=\mathcal{M}_{T_{V}}$.
Proof.
Let $\{a\prec b\prec c\prec d\}\subseteq V$. Then $\chi_{T}(a,b,c,d)=+1$ if and only if there is an arc $xy$ of $T$ with $a\prec b\preccurlyeq x\prec c\prec d\preccurlyeq y$. This then the case if and only if $a\prec b\preccurlyeq x^{V}\prec c\prec d\preccurlyeq y^{V}$. Finally, this is the case if and only if there exists an arc $x^{V}y^{V}$ of $T_{V}$ such that this holds, which is the condition for $\chi_{T_{V}}(a,b,c,d)=+1$. ∎
Hence, using [2, Corollary 3.6.3], in order to show that $\chi_{T}$ is a well-defined chirotope, it suffices to show that $\chi_{T}$ is realisable for $T$ a triangulation of a hexagon.
Proposition 3.7.
If $T$ is a hexagon triangulation, then $\chi_{T}$ is a realisable chirotope.
Proof.
If $T$ is the triangulation $\{26,36,46\}$, then one can check that $\chi_{T}$ is positive on all 4-subsets of $[6]$ ordered in the usual way. Hence, $\chi_{T}$ is realised by the three-dimensional cyclic polytope with six vertices [2, Section 9.4]. By rotating, we have that $\chi_{T}$ is realisable for any triangulation $T$ of the form $\{(a-4)a,(a-3)a,(a-2)a\}$.
Up to rotation, there are only three more hexagon triangulations we need to check, namely $T_{\Delta}$, $T_{\scriptsize\reflectbox{N}}$, and $T_{N}$ as shown in Figure 1. We claim that $\mathcal{M}_{T_{\Delta}}$, $\mathcal{M}_{T_{\tiny\reflectbox{N}}}$, and $\mathcal{M}_{T_{N}}$ are realised by the respective matrices
$M_{T_{\Delta}}=\left(\begin{smallmatrix}1&1&1&1&1&1\\ 0&1&2&3&4&5\\ 0&1&4&9&16&25\\ 3&1&-1&1&4&12\\ \end{smallmatrix}\right),\,M_{T_{\tiny\reflectbox{N}}}=\left(\begin{smallmatrix}1&1&1&1&1&1\\ 0&1&2&3&4&5\\ 0&1&4&9&16&25\\ 10&1&-1&1&11&32\end{smallmatrix}\right),\,M_{T_{N}}=\left(\begin{smallmatrix}1&1&1&1&1&1\\ 0&1&2&3&4&5\\ 0&1&4&9&16&25\\ 4&1&-1&1&11&26\end{smallmatrix}\right).$
Indeed, the maximal minors of these matrices are as shown in Table 1. We leave it to the reader to check that the signs of these minors agree with Definition 3.1. ∎
[…]
Figure 1. Three hexagon triangulations $T$ for which it suffices to check that $\mathcal{M}_{T}$ is realisable
Column tuple Minor of $M_{T_{\Delta}}$ Minor of $M_{T_{\tiny\reflectbox{N}}}$ Minor of $M_{T_{N}}$
1234 8 6-6 6
1235 18 10-10 26
1236 38 6-6 54
1245 6 6 42
1246 34 42 102
1256 48 64 64
1345 10-10 18 30
1346 6-6 72 72
1356 34 102 42
1456 38 54 6-6
2345 6-6 8 8
2346 10-10 30 18
2356 6 42 6
2456 18 26 10-10
3456 8 6 6-6
Table 1. Minors of the matrices $M_{T_{\Delta}}$, $M_{T_{\tiny\reflectbox{N}}}$, and $M_{T_{N}}$
Hence, we obtain the following result by applying [2, Corollary 3.6.3].
Theorem 3.8.
We have that $\chi_{T}$ is a well-defined chirotope, and so $\mathcal{M}_{T}$ is a well-defined oriented matroid.
It will be useful to describe the signed circuits of $\mathcal{M}_{T}$.
Proposition 3.9.
Every signed circuit of $\mathcal{M}_{T}$ is of the form $C=(\{a,c,e\},\{b,d\})$ or $C=(\{b,d\},\{a,c,e\})$ for $a\prec b\prec c\prec d\prec e$.
Proof.
First, note that every 5-subset of $[m]$ is a circuit of $\mathcal{M}_{T}$, since every 4-subset is a basis. Thus, let $\underline{C}=\{a\prec b\prec c\prec d\prec e\}\subseteq[m]$ be the underlying set of a signed circuit of $\mathcal{M}_{T}$, where $a\prec b\prec c\prec d\prec e$. Clearly the signs of $C$ are determined by the restriction $\mathcal{M}_{T}(\underline{C})=\mathcal{M}_{T_{\underline{C}}}$. We have that $T_{\underline{C}}$ is a triangulation of a pentagon. Up to rotation, there is only one triangulation of a pentagon, so we can rotate and assume that $T_{\underline{C}}=\{be,ce\}$. We then have that $\mathcal{M}_{T_{\underline{C}}}$ is then the oriented matroid of a three-dimensional cyclic polytope with five vertices. This implies that $C=(\{a,c,e\},\{b,d\})$ or $C=(\{b,d\},\{a,c,e\})$ by [3]. ∎
Corollary 3.10.
$\mathcal{M}_{T}$ is acyclic.
We can now describe the facets of the matroids $\mathcal{M}_{T}$. We write $T\pm 1$ for the triangulation of the $m$-gon with triangles $H\pm 1$ where $H$ is a triangle of $T$.
Proposition 3.11.
The facets of the oriented matroid $\mathcal{M}_{T}$ are given by $H$ where either $H$ is a triangle of $T$ or $H$ is a triangle of $T+1$.
We refer to facets $H$ where $H$ is a triangle of $T$ as upper facets of $\mathcal{M}_{T}$ and facets $H$ where $H$ is a triangle of $T+1$ as lower facets. This follows standard practice for cyclic polytopes; see [7]. In order to prove Proposition 3.11, we will need to use [2, Lemma 3.5.8], which we reproduce for the convenience of the reader.
Lemma 3.12 ([2, Lemma 3.5.8]).
Let $\chi$ be a chirotope on a set $E$ with a corresponding oriented matroid $\mathcal{M}$. Let $\underline{D}$ be the underlying set of a signed cocircuit of $\mathcal{M}$. Given $e,f\in\underline{D}$ with $e\neq f$, set
$s(e,f)=\chi(e,x_{2},\dots,x_{r})\cdot\chi(f,x_{2},\dots,x_{r})\in\{-1,+1\}$
where $X=(x_{2},\dots,x_{r})$ is an ordered basis of the hyperplane $E\setminus\underline{D}$ of $\mathcal{M}$. Then $s(e,f)$ does not depend upon the choice of $X$.
The signatures of $\underline{D}$ are given by $(D^{+},D^{-})$ and $(D^{-},D^{+})$, where
$D^{+}$ $=\{e\}\cup\{f\in\underline{D}\setminus e:s(e,f)=+1\},$
$D^{-}$ $=\{e\}\cup\{f\in\underline{D}\setminus e:s(e,f)=-1\}.$
Moreover, this pair of signatures does not depend upon the choice of $e$.
Proof of Proposition 3.11.
We first show that every hyperplane given by a triangle $H$ of $T$ is a facet. Let $H=abc$ be a triangle of $T$. We must show that every $x\in[m]\setminus H$ has the same sign in this cocircuit. If $c\prec x\prec a$, then $\chi_{T}(x,a,b,c)=+1$ due to the arc $ac$ of $T$. If $a\prec x\prec b$, then $\chi_{T}(x,a,b,c)=-\chi_{T}(a,x,b,c)=+1$ due to the arc $ab$ of $T$. One can similarly verify for $b\prec x\prec c$ that we have $\chi_{T}(x,a,b,c)=+1$. Using Lemma 3.12, we conclude that the signatures of $m]\setminus H$ are $([m]\setminus H,\varnothing)$ and $(\varnothing,[m]\setminus H)$, and so $H$ is a facet of $\mathcal{M}_{T}$.
Now let $H=abc$ be such that H1H-1 is a triangle of $T$. If $x\in[m]\setminus H$ has $c\prec x\prec a$, then $\chi_{T}(x,a,b,c)=-1$ due to the arc (a1)(c1)(a-1)(c-1) of $T$. Proceeding similarly to above, it is clear that $x$ must always have the sign 1-1 whenever it lies in $m]\setminus H$. We conclude that $H$ is a facet of $\mathcal{M}_{T}$ in this case too.
We now show that all facets of $\mathcal{M}_{T}$ are either upper facets or lower facets. Indeed, suppose that abcabc is a facet of $\mathcal{M}_{T}$. By rotating, we may assume that $c=m$ and $a\neq 1$. Suppose that $\chi(1,a,b,c)=+1$. Then there must be an arc $xc$ of $T$ with $a\preccurlyeq x\prec b$. If $x\neq a$, then $\chi(x,a,b,c)=-\chi(a,x,b,c)=-1$ due to the arc $xc$, which contradicts the fact that abcabc is a facet. We conclude that $ac$ is an arc of $T$. One can show in a similar way that $ab$ and $bc$ are arcs of $T$ if they are not boundary edges of the $m$-gon, so that abcabc is an upper facet. A similar argument shows that if $\chi(1,a,b,c)=-1$, then abcabc is a lower facet. Hence, all facets of $\mathcal{M}_{T}$ are either upper facets or lower facets. ∎
Given a basis $\sigma$ of $\mathcal{M}_{T}$, one can define upper and lower facets of $\sigma$ as the upper and lower facets of the matroid $\mathcal{M}_{T}(\sigma)$. Explicitly, we have the following.
Definition 3.13.
Let $\sigma=\{a,b,c,d\}$ be a basis of $\mathcal{M}_{T}$ with $a\prec b\prec c\prec d$. Then if $\chi_{T}(a,b,c,d)=+1$, we say that $\{a,b,d\}$ and $\{b,c,d\}$ are upper facets of $\sigma$ with respect to $T$ and $\{a,b,c\}$ and $\{a,c,d\}$ are lower facets of $\sigma$ with respect to $T$.
Lemma 3.14.
Given a triangulation $\Delta$ of $\mathcal{M}_{T}$ with $\sigma,\tau\in\Delta$, if $|\sigma\cap\tau|=3$, then $\sigma\cap\tau$ is an upper facet of one of $\sigma$ or $\tau$, and a lower facet of the other.
Proof.
Suppose for contradiction that $\sigma\cap\tau$ is a lower facet of both $\sigma$ and $\tau$, with the case where it is an upper facet of both being similar. By Proposition 3.6, it suffices to show the claim when restricting to $\sigma\cup\tau$. However, $|\sigma\cup\tau|=5$, whence the restriction $\mathcal{M}_{T}(\sigma\cup\tau)$ is the oriented matroid of three-dimensional cyclic polytope with five vertices by Proposition 3.9 and [3]. In this case, the notions of upper and lower facets coincide with the usual notions for simplices in cyclic polytopes; see, for instance, [7, Section 2]. For cyclic polytopes, it follows for geometric reasons that if $\sigma\cap\tau$ is a lower facet of both $\sigma$ and $\tau$ then their geometric realisations $\sigma||$ and $\tau||$ intersect in more than this facet, since points a small distance above $\sigma||$ and $\tau||$ must lie in both $\sigma||$ and $\tau||$. It follows from this that there must be a circuit $C=(C^{+},C^{-})$ such that $\sigma\supseteq C^{+}$ and $\tau\supseteq C^{-}$. In fact, since $|\underline{C}|=5=|\sigma\cup\tau|$ and $|\sigma|=|\tau|=4$, if we let $\sigma\setminus\tau=\{a\}$, then it follows that $a\in C^{+}$ and $\underline{C}\setminus\{a\}\subseteq\tau$. Hence, $\sigma$ and $\tau$ overlap on a circuit, meaning that they cannot be elements of the same triangulation, and so we have arrived at a contradiction. ∎
Noting Lemma 3.14, we make the following definition.
Definition 3.15.
Given a triangulation $\Delta$ of $\mathcal{M}_{T}$, we define a relation $\rightsquigarrow$ on $\Delta$ to be the smallest reflexive transitive relation such that $\sigma\rightsquigarrow\tau$ whenever $\sigma\cap\tau$ is an upper facet of $\sigma$ and a lower facet of $\tau$. Following [19, Definition 2.13], we say that a triangulation $\Delta$ of $\mathcal{M}_{T}$ is stackable if the relation $\rightsquigarrow$ is a partial order on $\Delta$.
In fact, we conjecture the following.
Conjecture 3.16.
Every triangulation $\Delta$ of $\mathcal{M}_{T}$ is stackable.
4. Showing the bijection
We now show that stackable triangulations of these oriented matroids are in bijection with equivalence classes of maximal green sequences
4.1. Preliminary results
We first show the following key result, which shows the relation between our oriented matroid $\mathcal{M}_{T}$ and the choice of extriangulated structure on $\mathscr{C}_{n}$ coming from the cluster tilting object $\mathsf{T}$ corresponding to $T+1$.
Lemma 4.1.
Let $\mathsf{T}\in\mathscr{C}_{n}$ be a cluster-tilting object, and let $T$ be the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}$. Let $\mathsf{O}_{A},\mathsf{O}_{B}\in\mathscr{C}_{n}$ be such that $A=\{a,c\}$, $B=\{b,d\}$ with $a\prec b\prec c\prec d$. Then $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})\neq 0$ if and only if $\chi_{T}(a,b,c,d)=+1$; and so $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})=0$ if and only if $\chi_{T}(a,b,c,d)=-1$.
Proof.
Note that by Section 2.2, we have that $\operatorname{Ext}_{\mathscr{C}_{n}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})\neq 0$ due to the fact that $a\prec b\prec c\prec d$. By definition of the extriangulated structure defined by $\mathsf{T}$, we have that $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})\neq 0$ if and only if the non-zero map $\mathsf{O}_{B}\to\Sigma\mathsf{O}_{A}$ factors through $\Sigma\mathsf{T}$. This is then the case precisely if there is a non-zero summand $\mathsf{O}_{X}$ of $\Sigma\mathsf{T}$ such that there is a non-zero composition $\mathsf{O}_{B}\to\mathsf{O}_{X}\to\Sigma\mathsf{O}_{A}$, since the map $\mathsf{O}_{B}\to\Sigma\mathsf{O}_{A}$ is unique up to scalar, by Section 2.2. By Lemma 4.2 and the description of $\Sigma$ from Section 2.2, this happens if and only if we have $b\preccurlyeq x\prec c\prec d\preccurlyeq y\prec a$, where $X=\{x,y\}$, and this is the case if and only if $\chi_{T}(a,b,c,d)=+1$. ∎
Lemma 4.2.
A pair of non-zero morphisms $\mathsf{O}_{P}\to\mathsf{O}_{Q}$ and $\mathsf{O}_{Q}\to\mathsf{O}_{R}$ in $\mathscr{C}_{n}$ compose to give a non-zero morphism $\mathsf{O}_{P}\to\mathsf{O}_{R}$ if and only if $p_{0}-1\preccurlyeq q_{0}-1\prec r_{0}\prec p_{1}-1\preccurlyeq q_{1}-1\prec r_{1}$, where $P=\{p_{0},p_{1}\}$, $Q=\{q_{0},q_{1}\}$, and $R=\{r_{0},r_{1}\}$.
Proof.
We argue using the bounded derived category $\mathscr{D}_{n}:=\mathscr{D}^{b}(K\mathbb{A}_{n})$. By [16, Lemma 6.6, Lemma 6.7, Proof of Proposition 6.1], the indecomposable objects of $\mathscr{D}_{n}$ may be labelled $\mathsf{U}_{i_{0}i_{1}}$ for $(i_{0},i_{1})\in\mathbb{Z}^{2}$ with $i_{0}+2\leqslant i_{1}$ and $i_{1}\leqslant i_{0}+n+1$, with $\operatorname{Hom}_{\mathscr{D}_{n}}(\mathsf{U}_{i_{0}i_{1}},\mathsf{U}_{j_{0}j_{1}})\neq 0$ if and only if $i_{0}-1<j_{0}<i_{1}-1<j_{1}<i_{0}+n+2$.
Suppose that $\mathsf{O}_{P}\to\mathsf{O}_{Q}$ and $\mathsf{O}_{Q}\to\mathsf{O}_{R}$ in $\mathscr{C}_{n}$ compose to give a non-zero morphism $\mathsf{O}_{P}\to\mathsf{O}_{R}$. Then, since $\mathscr{C}_{n}$ is an orbit category of $\mathscr{D}_{n}$, we can lift these morphisms to morphisms $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{Q^{\prime}}$ and $\mathsf{U}_{Q^{\prime}}\to\mathsf{U}_{R^{\prime}}$ in $\mathscr{D}_{n}$ which compose to give a non-zero map, where $\psi(\mathsf{U}_{P^{\prime}})=\mathsf{O}_{P}$, $\psi(\mathsf{U}_{Q^{\prime}})=\mathsf{O}_{Q}$, and $\psi(\mathsf{U}_{R^{\prime}})=\mathsf{O}_{R}$, with $\psi\colon\mathscr{D}_{n}\to\mathscr{C}_{n}$ the functor sending an object to its orbit. We have non-zero morphisms $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{Q^{\prime}}$ and $\mathsf{U}_{Q^{\prime}}\to\mathsf{U}_{R^{\prime}}$ in $\mathscr{D}_{n}$ composing to give a non-zero morphism $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{R^{\prime}}$ if and only if there are non-zero morphisms between these three pairs of objects, which is the case if and only if $p^{\prime}_{0}-1<q^{\prime}_{0}<p^{\prime}_{1}-1<q^{\prime}_{1}<p^{\prime}_{0}+n+2$, and $q^{\prime}_{0}-1<r^{\prime}_{0}<q^{\prime}_{1}-1<r^{\prime}_{1}<q^{\prime}_{0}+n+2$, and $p^{\prime}_{0}-1<r^{\prime}_{0}<p^{\prime}_{1}-1<r^{\prime}_{1}<p^{\prime}_{0}+n+2$. These three inequalities then all hold if and only if
(4.1) $p^{\prime}_{0}-1\leqslant q^{\prime}_{0}-1<r^{\prime}_{0}<p^{\prime}_{1}-1\leqslant q^{\prime}_{1}-1<r^{\prime}_{1}<p^{\prime}_{0}+n+2.$
By [16, Lemma 6.7], we have that $\psi(\mathsf{U}_{i^{\prime}_{0}i^{\prime}_{1}})=\mathsf{O}_{i_{0},i_{1}}$ where $\{i_{0},i_{1}\}=\{i^{\prime}_{0}\mod n+3,i^{\prime}_{1}\mod n+3\}$, and so (4.1) gives us the desired cyclic ordering $p_{0}-1\preccurlyeq q_{0}-1\prec r_{0}\prec p_{1}-1\preccurlyeq q_{1}-1\prec r_{1}$.
Conversely, if such a cyclic ordering holds, one can lift $\mathsf{O}_{P}$, $\mathsf{O}_{Q}$, and $\mathsf{O}_{R}$ to objects $\mathsf{U}_{P^{\prime}}$, $\mathsf{U}_{Q^{\prime}}$, and $\mathsf{U}_{R^{\prime}}$ of $\mathscr{D}_{n}$ such that (4.1) holds. This gives that the morphisms in $\mathscr{D}_{n}$ compose to give a non-zero morphism, and so the morphisms in $\mathscr{C}_{n}$ likewise compose to give a non-zero morphism. ∎
Definition 4.3.
Given a cluster-tilting object $\mathsf{T}$ in $\mathscr{C}_{n}$, we thus define $\mathcal{M}_{\mathsf{T}}$ to be the oriented matroid $\mathcal{M}_{T}$, where $T$ is the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}$. We similarly write $\chi_{\mathsf{T}}=\chi_{T}$.
4.2. Maximal green sequence to triangulation
We explain how to construct a triangulation of $\mathcal{M}_{\mathsf{T}}$ from a maximal green sequence of $\mathsf{T}$.
Construction 4.4.
Let $\mathcal{G}$ be a maximal green sequence of $\mathsf{T}$, given by a sequence of cluster-tilting objects $\mathsf{T}=\mathsf{T}_{0},\mathsf{T}_{1},\dots,\mathsf{T}_{l}=\Sigma\mathsf{T}$ in $\mathscr{C}_{n}$. Denote the exchange pair of the mutation from $\mathsf{T}_{i-1}$ to $\mathsf{T}_{i}$ by $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$. Since we have $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{B_{i}},\mathsf{O}_{A_{i}})\neq 0$, the arcs defined by $A_{i}$ and $B_{i}$ cross in the $(n+3)$-gon, and so $|A_{i}\cup B_{i}|=4$. Hence, we define the set
$\Delta_{\mathcal{G}}:=\{A_{i}\cup B_{i}:(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})\in\mathrm{Exch}(\mathcal{G})\}$
of $3$-simplices of $\mathcal{M}_{\mathsf{T}}$. We claim that this is indeed a triangulation of $\mathcal{M}_{\mathsf{T}}$.
Lemma 4.5.
If $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are equivalent, then $\Delta_{\mathcal{G}}=\Delta_{\mathcal{G}^{\prime}}$.
Proof.
If $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are equivalent, then $\mathrm{Exch}(\mathcal{G})=\mathrm{Exch}(\mathcal{G}^{\prime})$, so $\Delta_{\mathcal{G}}=\Delta_{\mathcal{G}^{\prime}}$. ∎
Proposition 4.6.
We have that $\Delta_{\mathcal{G}}$ is a triangulation of $\mathcal{M}_{\mathsf{T}}$.
Proof.
We verify the properties of Definition 2.5. For (1), suppose that $\sigma,\tau\in\Delta_{\mathcal{G}}$. We have that $\mathcal{M}_{\mathsf{T}}(\sigma)$ and $\mathcal{M}_{\mathsf{T}}(\tau)$ are both the uniform rank-four oriented matroid on a 4-element set. Hence, all subsets of size three of $\sigma$ and $\tau$ are facets of these respective matroids and all subsets of $\sigma$ and $\tau$ are faces of $\mathcal{M}_{\mathsf{T}}(\sigma)$ and $\mathcal{M}_{\mathsf{T}}(\tau)$ respectively. It follows that $\sigma\cap\tau$ must be a common face of $\mathcal{M}_{\mathsf{T}}(\sigma)$ and $\mathcal{M}_{\mathsf{T}}(\tau)$.
We now verify that (2) holds. Let $T$ be the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}$. Suppose that $\sigma\in\Delta_{\mathcal{G}}$. Indeed, let $\sigma=A_{i}\cup B_{i}$. Let $F$ be a facet of $\sigma$ containing $A_{i}$; one can argue similarly for facets of $\sigma$ containing $B_{i}$. If $F$ is a triangle of $T+1$, then it is a facet of $\mathcal{M}_{T}$ by Proposition 3.11.
Thus, suppose that $F$ is not a triangle of $T+1$. Then, if $T_{j}$ is the triangulation corresponding to $\mathsf{T}_{j}$, let $T_{k}$ be the first triangulation where $F$ appears. By assumption, $k\neq 0$. There is then an exchange pair $(\mathsf{O}_{A_{k}},\mathsf{O}_{B_{k}})$ and a simplex $\sigma^{\prime}=A_{k}\cup B_{k}\in\Delta_{\mathcal{G}}$. Since $F$ is a triangle of $T_{k}$ but not of $T_{k-1}$, we have that $F$ must also be a facet of $\sigma^{\prime}$, and must contain $B_{k}$. This means that $\sigma\neq\sigma^{\prime}$, and so $F$ is contained in at least two simplices of $\Delta_{\mathcal{G}}$.
We must now show that $F$ cannot be contained in more than two simplices of $\Delta_{\mathcal{G}}$. Suppose that we have another simplex $\sigma^{\prime\prime}\in\Delta_{\mathcal{G}}$ with $F$ as a facet of $\sigma^{\prime\prime}$. We have a total order on $\Delta_{\mathcal{G}}$ given by the order of the exchange pairs of $\mathcal{G}$. When a simplex $\tau$ corresponds to the exchange pair from $\mathsf{T}_{h-1}$ to $\mathsf{T}_{h}$, the lower facets of $\tau$ leave $T_{h-1}$ and are replaced by the upper facets of $\tau$. Since we have at least three simplices in $\Delta_{\mathcal{G}}$ which contain $F$, we have that $F$ enters the sequence of triangulations $T_{0},T_{1},\dots,T_{l}$, leaves the sequence, and then enters again. In particular, this must be true for some $1$-dimensional face $A$ of $F$, which we can choose to be an arc in the $(n+3)$-gon, since $n+3\geqslant 4$. This then gives us an indecomposable object $\mathsf{O}_{A}$ which enters $\mathcal{G}$, leaves, and then enters again. But this contradicts [11, Lemma 3.4].
Finally, we must show (3): that no two simplices $\sigma,\tau\in\Delta$ overlap on a circuit. Suppose for contradiction that, in fact, there is a circuit $(C^{+},C^{-})$ of $\mathcal{M}_{T}$ such that $\sigma\supseteq C^{+}$ and $\tau\supseteq\underline{C}\setminus\{x\}$ for some $x\in C^{+}$. In particular, $\tau\supseteq C^{-}$. By rotating, we can assume that $\underline{C}=\{a\prec b\prec c\prec d\prec e\}$ with $C^{\pm}=\{a,c,e\}$ and $C^{\mp}=\{b,d\}$ by Proposition 3.9. Without loss of generality, suppose that $C^{+}=\{a,c,e\}$ and $C^{-}=\{b,d\}$. Suppose now that $\sigma$ precedes $\tau$ in the order given by $\mathcal{G}$. We have that $\chi_{\mathsf{T}}(b,c,d,e)=+1$ since the description of $C$ gives us that $\mathcal{M}_{T}(\underline{C})$ is the oriented matroid of a three-dimensional cyclic polytope with five vertices. Hence $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{ce},\mathsf{O}_{bd})\neq 0$ by Lemma 4.1. However, since $\sigma$ precedes $\tau$, we have that $\mathsf{O}_{ce}$ must precede $\mathsf{O}_{bd}$ in $\mathcal{G}$, which contradicts [11, Lemma 3.1(2)]. If $\tau$ precedes $\sigma$, then note that $\chi_{T}(a,b,c,d)=+1$, so $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{bd},\mathsf{O}_{ac})\neq 0$, which contradicts [11, Lemma 3.1(2)] in a similar way. ∎
Proposition 4.7.
The triangulation $\Delta_{\mathcal{G}}$ is stackable.
Proof.
First note that $\mathcal{G}$ gives a total order $\leqslant_{\mathcal{G}}$ on its exchange pairs according to the order they appear. Since the exchange pairs of $\mathcal{G}$ are in bijection with the simplices of $\Delta_{\mathcal{G}}$, we have that $\leqslant_{\mathcal{G}}$ is also a total order on the simplices of $\Delta$. We claim that the relation $\rightsquigarrow$ on $\Delta$ is contained in the relation $\leqslant_{\mathcal{G}}$. From this it follows that $\rightsquigarrow$ is a partial order, since any reflexive transitive relation contained in a total order is a partial order.
In order to show this, it suffices to show that whenever we have two simplices $\sigma$ and $\tau$ of $\Delta$ such that $\sigma\cap\tau$ is an upper facet of $\sigma$ and a lower facet of $\tau$, we have that $\sigma\leqslant_{\mathcal{G}}\tau$. Let $A_{i}$ be the intersection of the lower facets of $\sigma$, with $B_{i}$ the intersection of the upper facets; similarly, let $A_{j}$ be the intersection of the lower facets of $\tau$, with $B_{j}$ the intersection of the upper facets.
By Definition 3.13 and Lemma 4.1, we have that $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ and $(\mathsf{O}_{A_{j}},\mathsf{O}_{B_{j}})$ are the respective exchange pairs corresponding to $\sigma$ and $\tau$. Let $\mathsf{T}_{j-1}$ be the cluster-tilting object of $\mathcal{G}$ on which the mutation of $\mathsf{O}_{A_{j}}$ for $\mathsf{O}_{B_{j}}$ is performed. Then, if $T_{j-1}$ is the triangulation of the $(n+3)$-gon corresponding to $\mathsf{T}_{j-1}$, we have that $T_{j-1}$ contains the lower facets of $\tau$. Since one of the lower facets of $\tau$ is an upper facet of $\sigma$, and each upper facet of $\sigma$ contains $B_{i}$, we have that $T_{j-1}$ contains the arc $B_{i}$. Then [11, Lemma 3.4] ensures that $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ occurs before $\mathsf{T}_{j-1}$, and so $\sigma\leqslant_{\mathcal{G}}\tau$. ∎
4.3. Triangulation to maximal green sequence
We now describe the inverse construction, which starts with a stackable triangulation $\Delta$ of $\mathcal{M}_{\mathsf{T}}$, and outputs an equivalence class of maximal green sequences $\mathcal{G}]$ of $\mathsf{T}$. As in the previous section, we have fixed a cluster-tilting object $\mathsf{T}$ in $\mathscr{C}_{n}$, and have let $T$ be the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}$. Given a triangulation $T_{i}$ of the $m$-gon which contains a quadrilateral abcdabcd, then flipping the diagonal in abcdabcd is called the increasing flip given by abcdabcd if it replaces the lower facets of abcdabcd with respect to $T_{i}$ by the upper facets of abcdabcd with respect to $T_{i}$. If the reverse is true, then it is called a decreasing flip.
Proposition 4.8.
Let $\Delta$ be a stackable triangulation of $\mathcal{M}_{T}$ with $\leqslant$ a linear extension of the partial order $\rightsquigarrow$ on the simplices of $\Delta$. Letting the total order $\leqslant$ on the simplices of $\Delta$ be $\sigma_{1}<\sigma_{2}<\dots<\sigma_{l}$, there is then a sequence of triangulations $T_{0},T_{1},\dots,T_{l}$ of the $(n+3)$-gon such that (1) $T_{0}=T+1$; (2) $T_{i}$is the increasing flip of $T_{i-1}$ given by $\sigma_{i}$; (3) $T_{l}=T$.
Proof.
We construct the sequence $T_{0},T_{1},\dots,T_{l}$ inductively by letting $T_{0}=T$ and letting $T_{i}$ be the increasing flip of $T_{i-1}$ given by $\sigma_{i}$. We show that this is well-defined, that is: the lower facets of $\sigma_{i}$ are in fact contained in $T_{i-1}$.
For the base case, $i=1$, let $F$ be a lower facet of $\sigma_{1}$. By Definition 2.5(2), we have that $F$ is either a facet of $\mathcal{M}_{T}$, or is contained in precisely two simplices of $\Delta$. If $F$ is contained in another simplex $\sigma_{j}$ of $\Delta$, then it must be an upper facet of $\sigma_{j}$ by Lemma 3.14. However, then we would have $\sigma_{j}\rightsquigarrow\sigma_{1}$, which would imply that $\sigma_{j}<\sigma_{1}$, since $\leqslant$ is a linear extension of $\rightsquigarrow$. Hence, all lower facets of $\sigma_{1}$ are facets of $\mathcal{M}_{T}$, and hence lower facets of $\mathcal{M}_{T}$. Since the lower facets of $\mathcal{M}_{T}$ are given by $T+1$, the base case holds.
Now, for the inductive step, let $F$ be one of the lower facets of $\sigma_{i}$. Again, we have that $F$ is either a facet of $\mathcal{M}_{T}$, or is contained in precisely two simplices of $\Delta$.
Suppose first that $F$ is a facet of $\mathcal{M}_{T}$. Then it must be a lower facet of $\mathcal{M}_{T}$. Thus, $F$ is a triangle of $T+1$. Moreover, $F$ cannot be contained in any other simplices of $\Delta$; in particular, this would contradict Lemma 3.14. Hence, $F$ must still be a triangle of $T_{i-1}$, since it could not have been removed by any other simplex.
Now suppose that $F$ is contained in precisely one other simplex $\sigma_{j}$ of $\Delta$. By Lemma 3.14, we have that $F$ must be an upper facet of $\sigma_{j}$, and so we have $\sigma_{j}\leqslant\sigma_{i}$. By the induction hypothesis, we have that the lower facets of $\sigma_{j}$ are contained in $T_{j-1}$, and so $T_{j}$ contains the upper facets of $\sigma_{j}$, and so contains $F$. Since $F$ is only contained in $\sigma_{j}$ and $\sigma_{i}$ by Definition 2.5(2), it remains a triangle of $T_{i-1}$.
We finally must show that (3) holds. Indeed, every upper facet of $\mathcal{M}_{T}$ must be an upper facet of some $\sigma_{i}$ by Definition 2.5(2) and must be an upper facet of exactly one $\sigma_{i}$ by Lemma 3.14. Hence, for every upper facet $F$ of $\mathcal{M}_{T}$, there must be some triangulation $T_{i-1}$ which contains the lower facets of the simplex $\sigma_{i}$, so that $T_{i}$ contains the upper facets of the simplex $\sigma_{i}$, and so contains $F$. After $\sigma_{i}$, there is no further $\sigma_{j}$ with $F$ as a facet, and so $T_{l}$ contains $F$. Therefore $T_{l}$ contains all upper facets of $\mathcal{M}_{T}$, and so $T_{l}=T$, as desired. ∎
Construction 4.9.
Let $\Delta$ be a stackable triangulation of $\mathcal{M}_{T}$. As before, since $\rightsquigarrow$ is a partial order on $\Delta$, we may take a linear extension of it to a total order $\leqslant$. Let $\sigma_{1},\sigma_{2},\dots,\sigma_{l}$ be the simplices of $\Delta$, ordered according to $\leqslant$.
We then take the sequence $T_{0}$, $T_{1}$, …, $T_{l}$ of triangulations from Proposition 4.8. By Section 2.2 and Lemma 4.1, this corresponds to a sequence of cluster-tilting objects $\mathsf{T}_{0}$, $\mathsf{T}_{1}$, …, $\mathsf{T}_{l}$ of $\mathscr{C}_{n}$ such that $\mathsf{T}_{0}=\mathsf{T}$, $\mathsf{T}_{i}$ is a green mutation of $\mathsf{T}_{i-1}$, and $\mathsf{T}_{l}=\Sigma\mathsf{T}$. Thus this gives a maximal green sequence $\mathcal{G}^{\leqslant}_{\Delta}$ of $\mathsf{T}$, and hence an equivalence class $\mathcal{G}^{\leqslant}_{\Delta}]$.
Lemma 4.10.
The equivalence class of maximal green sequences $\mathcal{G}^{\leqslant}_{\Delta}]$ is independent of the linear extension $\leqslant$ chosen of the partial order $\rightsquigarrow$.
Proof.
If $\leqslant^{\prime}$ is a different linear extension, then we have $\mathrm{Exch}(\mathcal{G}^{\leqslant}_{\Delta})=\mathrm{Exch}(\mathcal{G}^{\leqslant^{\prime}}_{\Delta})$, since both correspond to the set of exchange pairs $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ where $\sigma_{i}=A_{i}\cup B_{i}$ is a simplex of $\Delta$, with $A_{i}$ the intersection of its lower facets and $B_{i}$ the intersection of its upper facets. In turn, this is because, for either $\leqslant$ or $\leqslant^{\prime}$, if $T_{i}$ is the increasing flip of $T_{i-1}$ given by $\sigma_{i}$, then $T_{i}$ is the result of replacing the arc $A_{i}$ of $T_{i-1}$ with the arc $B_{i}$. By the relation between triangulations of the $(n+3)$-gon and cluster-tilting objects from Section 2.2, we then have that the corresponding cluster-tilting objects $\mathsf{T}_{i}$ and $\mathsf{T}_{i-1}$ are related by the mutation with exchange pair $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$. ∎
Since our constructions in this subsection and the previous are inverse to each other, we obtain the following theorem.
Theorem 4.11.
Let $\mathsf{T}$ be a basic cluster-tilting object in $\mathscr{C}_{n}$. There is then a bijection between stackable triangulations of $\mathcal{M}_{\mathsf{T}}$ and equivalence classes of maximal green sequences of $\mathsf{T}$.